Hi I need a little refresher in point-slope and slope intercepter and need the full steps with Slope = –5, passing through (–4, –2) and Slope = –2, passing through (0, –3)
Thank you

Answer by scott8148(6628)

slope-intercept ___ y = mx + b ___ m is slope, b is y-intercept

plug in slope, then plug in coordinates to find b

I'll do the 1st one ___ y = -5x + b ___ -2 = -5(-4) + b ___ 18 = b ___ y = -5x + 18